# Back to the origins of brane-antibrane inflation

### v1

@@ 4.2 Case without linear contribution
Figure 4: The red curve is the inflationary potential for the case without linear contribution which features a dS minimum with $V_{\rm min}^{\rm inf}\sim 10^{-13}$, while the blue curve is the late-time potential with a much smaller vacuum energy but with a minimum at similar field values. The parameters used to generate this plot are $\xi=g_{s}=0.1$, $W_{0}=1$, $W_{X}^{2}=10^{-7}$ and $T_{7}=2\pi$, while $\alpha$ has been fixed as in (35) with $\theta=0.994$.
@@ 5 Allowed UV parameter space
  • Compatibility with data: We use the Planck data Planck:2018jri for observational constraints on the amplitude of density perturbations and the spectral index $n_{s}$. It can be shown Burgess:2022nbx that CMB observations can be matched at horizon exit around $N_{e}\simeq 56$ e-foldings before the end of inflation where:
    $V_{\rm inf}\simeq 10^{-17}\,M_{p}^{4}\,,\qquad\varphi_{*}\simeq 10^{-3}M_{p}\qquad\text{and}\qquad r\simeq 2\times 10^{-8}\,,$ (55)
@@
      $\frac{m_{3/2}}{M_{s,{\rm warped}}}\simeq\frac{g_{s}^{1/4}W_{0}}{\sqrt{2}\,\sigma_{\rm min}}\,e^{\rho}\ll 1\qquad\Leftrightarrow\qquad\sigma_{\rm min}\gg\frac{g_{s}^{1/4}W_{0}}{\sqrt{2}}\,e^{\rho}\,.$ (57)
@@
      Dilute flux approximation under control: Large throats at not-too-large-volumes can lead to a failure of the dilute flux approximation, thus modifying the computation of the 4D Planck mass and yielding corrections to the Einstein-Hilbert term Gao:2022fdi . Even though this does not need to be problematic for our proposed inflation mechanism (since the modified corrections could still be used for uplifting to the correct vacuum energy), we will be conservative and impose:
@@
With linear contribution $10$ $0.15$ $\simeq 0.24$ $29$ $\simeq 3360$
Without linear contribution $10$ $0.15$ $\simeq 0.16$ $23$ $\simeq 670$
@@
Albeit a bit large, the tadpoles arising from the flux integers shown in Tab. 1, $KM=290$ and $KM=230$ for the two cases respectively, could in principle be cancelled in realistic models Kreuzer:2000xy .
@@ 6 Conclusions
In this paper we derived the scalar potential of brane-antibrane inflation within a low-energy EFT of type IIB string compactifications where the volume mode is stabilised by perturbative effects and inflation is driven by the motion of a D3-brane that is attracted by an $\overline{\rm D3}$-brane at the tip of a warped throat. The key features of our construction are the following ones:
@@
Let us stress that our new framework for brane-antibrane inflation based on perturbative moduli stabilisation avoids the problems faced by the original proposal but shares its interesting aspects. In particular, the end of inflation and reheating can have a purely stringy nature, contrary to many other string inflationary models where this epoch can instead be described within a low-energy supergravity perspective. In fact, brane-antibrane inflation typically ends as in hybrid inflation scenarios with the open string inflaton that becomes the tachyonic waterfall field at a critical interbrane separation. The development of a tachyon captures the process of brane-antibrane annihilation Sen:1998sm ; Sen:2002nu ; Sen:2002in whose endpoint is a gas of highly excited strings. These string modes eventually decay giving rise to successful reheating Kofman:2005yz . Before decaying, these strings reach thermal equilibrium (in the presence of branes) in a Hagedorn phase Frey:2005jk ; Frey:2023khe which leads to the emission of gravitational waves peaked in the GHz band Frey:2024jqy (see Aggarwal:2020olq for a review). If the strings are instead stable, they behave as stringy cosmic strings Sarangi:2002yt ; Copeland:2003bj ; Majumdar:2005qc ; Conlon:2024uob that may be tested experimentally, independently of CMB tests. Additionally, hybrid inflation models have been found to generate primordial black holes (see for instance LISACosmologyWorkingGroup:2023njw ). Studying in detail the dynamics or reheating and exploring the possibility to generate primordial black holes is very interesting but it is beyond the scope of our paper, and so we leave it for further studies.
@@ Appendix A Brane-antibrane inflation in warped compactifications
$ds^{2}=\left(1+\frac{e^{-4A(z)}}{\mathcal{V}^{2/3}}\right)^{-1/2}\eta_{\mu\nu}dx^{\mu}dx^{\nu}+\left(1+\frac{e^{-4A(z)}}{\mathcal{V}^{2/3}}\right)^{1/2}g_{\alpha\bar{\beta}}dz^{\alpha}\,d\bar{z}^{\bar{\beta}}\,.$ (61)
@@
where $\gamma_{ab}$ is the pullback of the ambient space metric, $B_{ab}$ is the pullback of the Neveu-Schwarz two-form, and $F_{ab}$ are gauge fluxes on the worldvolume. Inflation is achieved due to the attractive nature of the interbrane potential that appears when one considers the backreaction of a brane in the antibrane action. Before introducing backreaction, it is useful to consider the dynamics of branes/antibranes in a warped background. To do so, let us consider a setup with no $B$-field or worldvolume fluxes. Taking an expansion in powers of derivatives and keeping only renormalisable terms, we find:
@@
$=$ $\int{d^{4}x\,\left(T_{3}\left(1+\frac{e^{-4A(r)}}{\mathcal{V}^{2/3}}\right)^{-1}+\frac{1}{2}g_{\alpha\bar{\alpha}}\partial_{\mu}(\sqrt{T_{3}}z^{\alpha})\partial^{\mu}(\sqrt{T_{3}}\bar{z}^{\bar{\alpha}})+\mathcal{O}(\partial z)^{4}\right)},$
@@
$k(z_{\alpha},\bar{z}_{\bar{\alpha}})=\frac{3}{2}\left(\sum_{A=1}^{4}|z_{A}|^{2}\right)^{2/3}=\frac{3}{2}\left(\sum_{\alpha=1}^{3}|z_{\alpha}|^{2}+\left|\sum_{\alpha=1}^{3}z_{\alpha}^{2}\right|\right)^{2/3}\,.$ (73)
@@
$1+\frac{e^{-4A}}{\mathcal{V}^{2/3}}\to 1+\frac{e^{-4A}}{\mathcal{V}^{2/3}}+\delta h(r)\,,\qquad-\nabla^{2}_{\tilde{r}}\delta h(\tilde{r};r)=\left(\frac{2\pi}{T_{3}}\right)\frac{\delta^{(6)}(\tilde{r}-r)}{\sqrt{g(\tilde{r})}}\,.$ (76)
@@ Appendix B Late-time potential with subleading terms
$K\simeq-3\ln\left[\tau-\alpha\ln\tau+\frac{\xi}{3cg_{s}^{3/2}\sqrt{\tau}}\left(c-g_{s}^{2}\ln\tau\left(1-\frac{1}{3\pi\ln\tau}\right)\right)\right],$ (79)
@@
$\frac{V}{3W_{0}^{2}}\simeq\frac{\alpha}{\tau^{4}}-\frac{\xi\sqrt{g_{s}}}{4c\tau^{9/2}}\left(\ln\tau-\frac{c}{g_{s}^{2}}-b\right)+\mathcal{O}\left(\frac{1}{\tau^{5}}\right)\quad\text{with}\quad b\equiv\frac{8}{3}+\frac{1}{3\pi}\simeq 2.8\,.$ (80)
@@
$\tau_{\rm min,new}=\lambda_{0}\,e^{b}\,e^{\frac{c}{g_{s}^{2}}}=e^{b}\,\tau_{\rm min}\quad\text{with}\quad\theta_{\rm new}\equiv\left(\frac{16c\,e^{\frac{10}{9}}}{9\xi\sqrt{g_{s}}}\right)\alpha\,e^{\frac{b}{2}}\,e^{\frac{c}{2g_{s}^{2}}}=e^{\frac{b}{2}}\,\theta\,.$ (81)
@@ C.1 Case with linear contribution
where:
$\mathcal{A}$ $=$ $\frac{1}{3}+\frac{g_{s}^{2}}{3\sigma}\left(1-2\ln\sigma+3(\ln\sigma)^{2}+\frac{2\alpha}{g_{s}^{2}}\ln\sigma\right)+...$
$\mathcal{B}$ $=$ $-2g_{s}-\frac{2g_{s}^{3}}{\sigma}\left[1-2\ln\sigma+3(\ln\sigma)^{2}+\frac{\alpha}{g_{s}^{2}}\left(\ln\sigma+2\right)\right]+...$
$\mathcal{C}$ $=$ $3\left(g_{s}^{2}+\alpha\right)+\frac{3\xi}{4g_{s}^{3/2}\sqrt{\sigma}}\left[1-\frac{g_{s}^{2}}{c}\left(\ln\sigma-b\right)\right]$ (83)
$+$ $\frac{3g_{s}^{4}}{\sigma}\left[3(\ln\sigma)^{2}+1+\frac{4\alpha}{g_{s}^{2}}+\frac{1}{g_{s}^{4}}\left(2\ln\sigma\left(\alpha^{4}-g_{s}^{4}\right)+3\alpha^{2}\right)\right]+...$
@@
$V_{\scriptscriptstyle NNLO}$ $\simeq$ $\left(\frac{W_{X}}{W_{0}}\right)^{a}\frac{W_{0}^{2}}{\sigma^{b}}\;\qquad\text{with}\qquad a+b=5\,.$
@@ Appendix D RG-induced moduli stabilisation and inflation
In this appendix we review the RG-inspired moduli stabilisation scenario proposed in Burgess:2022nbx with in mind the goal of deepening the analysis of brane-antibrane inflation. Let us consider a string-inspired supersymmetric 4D EFT where the superpotential $W$ does not depend on the Kähler modulus $T$, i.e. $W=W_{0}$, and the Kähler potential can be expanded as:
@@
$W=W_{0}(\Phi)+XW_{X}(\Phi,\overline{\Phi})\qquad\text{and}\qquad\sigma\equiv e^{-K/3}=\tau-\hat{k}+...\,,$ (89)
@@
$\hat{k}=\kappa(\Phi,\overline{\Phi},\ln\tau)+(X+\overline{X})\kappa_{X}(\Phi,\overline{\Phi},\ln\tau)+X\overline{X}\kappa_{X\overline{X}}(\Phi,\overline{\Phi},\ln\tau)\,.$ (90)
@@
$\kappa^{\overline{X}X}\sim\mathcal{O}(1)\qquad\kappa_{X\overline{T}}\sim\mathcal{O}\left(\frac{\alpha_{g,0}^{2}}{\sigma}\right)\qquad\kappa_{T\overline{T}}\sim\mathcal{O}\left(\frac{k_{1}\alpha_{g,0}^{2}}{\sigma^{2}}\right)$ (92)

### v2

@@ 4.2 Case without linear contribution
Figure 4: The red curve is the inflationary potential for the case without linear contribution which features a dS minimum with $V_{\rm min}^{\rm inf}\simeq 2.5\times 10^{-13}$, while the blue curve is the late-time potential with a much smaller vacuum energy but with a minimum at similar field values. The parameters used to generate this plot are $\xi=g_{s}=0.1$, $W_{0}=1$, $W_{X}^{2}=10^{-7}$ and $T_{7}=2\pi$, while $\alpha$ has been fixed as in (39) with $\theta=0.994$.
@@ 5 Allowed UV parameter space
  • Compatibility with data: We use the Planck data Planck:2018jri for observational constraints on the amplitude of density perturbations and the spectral index $n_{s}$. It can be shown Burgess:2022nbx that CMB observations can be matched at horizon exit around $N_{e}\simeq 56$ e-foldings before the end of inflation where $\varphi_{*}\simeq 10^{-3}M_{p}$ and:
    $V_{\rm inf}\simeq 10^{-17}\,M_{p}^{4}\,,\quad n_{s}\simeq 0.970\quad\text{and}\quad r\simeq 2\times 10^{-8}\,,$ (60)
@@
      $\frac{m_{3/2}}{M_{s,{\rm warped}}}\simeq\frac{g_{s}^{1/4}W_{0}}{\sqrt{2}\,\sigma_{\rm min}}\,e^{\rho}\ll 1\,,$ (62)
      which implies:
      $\sigma_{\rm min}\gg\frac{g_{s}^{1/4}W_{0}}{\sqrt{2}}\,e^{\rho}\,.$ (63)
@@
      Dilute flux approximation under control: Very large throats at not-too-large-volumes can lead to a failure of the dilute flux approximation, thus modifying the computation of the 4D Planck mass and yielding corrections to the Einstein-Hilbert term Gao:2022fdi . Even though this does not need to be problematic for our proposed inflation mechanism (since the modified corrections could still be used for uplifting to the correct vacuum energy), we will be conservative and impose:
@@
With linear term $10$ $0.15$ $\simeq 0.24$ $29$ $\simeq 3360$
Without linear term $10$ $0.15$ $\simeq 0.16$ $23$ $\simeq 670$
@@
Albeit a bit large, the tadpoles arising from the flux integers shown in Tab. 1, $KM=290$ and $KM=230$ for the two cases respectively, could in principle be cancelled in realistic models Kreuzer:2000xy . Let us finally point out that higher order $\alpha^{\prime}$ corrections might require a very large value of $g_{s}M^{2}$ to avoid the KPV decay of the $\overline{\rm D3}$-brane Schreyer:2022len ; Schreyer:2024pml . Even if the corresponding decay channel is not well understood for a single $\overline{\rm D3}$-brane, it is easy to check that imposing $g_{s}M^{2}\geq 150$, instead of $g_{s}M^{2}\geq 46$ as in (66), there is still a large allowed region of underlying parameter space. For instance, in the linear case, $g_{s}=0.05$, $K=10$ and $W_{0}=0.1$ would give $M=70$, and so $g_{s}M^{2}=245$. The only potential problem for this case would be a rather large tadpole, $KM=700$. We leave however a concrete realisation of our model for future work.
@@ 6 Conclusions
In this paper, expanding on a proposal in Burgess:2022nbx , we derived the scalar potential of brane-antibrane inflation within a low-energy EFT of type IIB string compactifications where the volume mode is stabilised by perturbative effects and inflation is driven by the motion of a D3-brane that is attracted by an $\overline{\rm D3}$-brane at the tip of a warped throat. The key features of our construction are the following ones:
@@
Let us stress that our new framework for brane-antibrane inflation based on perturbative moduli stabilisation avoids the problems faced by the original proposal but shares its interesting aspects. In particular, the end of inflation and reheating can have a purely stringy nature, contrary to many other string inflationary models where this epoch can instead be described within a low-energy supergravity perspective. In fact, brane-antibrane inflation typically ends as in hybrid inflation scenarios with the open string inflaton that becomes the tachyonic waterfall field at a critical interbrane separation. The arising of a tachyon Sen:1998sm ; Sen:2002nu ; Sen:2002in captures the process of brane-antibrane annihilation whose endpoint is a gas of highly excited strings. These string modes eventually decay giving rise to successful reheating Kofman:2005yz . Before decaying, these strings reach thermal equilibrium (in the presence of branes) in a Hagedorn phase Frey:2005jk ; Frey:2023khe which leads to the emission of gravitational waves peaked in the GHz band Frey:2024jqy (see Aggarwal:2020olq for a review). If the strings are instead stable, they behave as stringy cosmic strings Sarangi:2002yt ; Copeland:2003bj ; Majumdar:2005qc ; Conlon:2024uob that may be tested experimentally, independently of CMB tests. Additionally, hybrid inflation models have been found to generate primordial black holes (see for instance LISACosmologyWorkingGroup:2023njw ). Studying in detail the dynamics or reheating and exploring the possibility to generate primordial black holes is very interesting but it is beyond the scope of our paper, and so we leave it for further studies.
@@ Appendix A Brane-antibrane inflation in warped compactifications
$ds^{2}$ $=$ $\left(1+\frac{e^{-4A(z)}}{\mathcal{V}^{2/3}}\right)^{-1/2}\eta_{\mu\nu}dx^{\mu}dx^{\nu}$ (67)
$+$ $\left(1+\frac{e^{-4A(z)}}{\mathcal{V}^{2/3}}\right)^{1/2}g_{\alpha\bar{\beta}}dz^{\alpha}\,d\bar{z}^{\bar{\beta}}\,.$
@@
where $\gamma_{ab}$ is the pullback of the ambient space metric, $B_{ab}$ is the pullback of the Neveu-Schwarz two-form, and $F_{ab}$ are gauge fluxes on the worldvolume. Inflation is achieved due to the attractive nature of the interbrane potential that appears when one considers the backreaction of a brane in the antibrane action. Before introducing backreaction, we consider the dynamics of branes/antibranes in a warped background. To do so, let us consider a setup with no $B$-field or worldvolume fluxes. Taking an expansion in powers of derivatives and keeping only renormalisable terms, we find:
@@
$=$ $T_{3}\int{d^{4}x\,\left(1+\frac{e^{-4A(r)}}{\mathcal{V}^{2/3}}\right)^{-1}}$
$+$ $\int{d^{4}x\,\left(\frac{1}{2}g_{\alpha\bar{\alpha}}\partial_{\mu}(\sqrt{T_{3}}z^{\alpha})\partial^{\mu}(\sqrt{T_{3}}\bar{z}^{\bar{\alpha}})+\mathcal{O}(\partial z)^{4}\right)},$
@@
$k(z_{\alpha},\bar{z}_{\bar{\alpha}})$ $=$ $\frac{3}{2}\left(\sum_{A=1}^{4}|z_{A}|^{2}\right)^{2/3}$ (79)
$=$ $\frac{3}{2}\left(\sum_{\alpha=1}^{3}|z_{\alpha}|^{2}+\left|\sum_{\alpha=1}^{3}z_{\alpha}^{2}\right|\right)^{2/3}\,.$
@@
$1+\frac{e^{-4A}}{\mathcal{V}^{2/3}}\to 1+\frac{e^{-4A}}{\mathcal{V}^{2/3}}+\delta h(r)\,,$ (82)
with:
$-\nabla^{2}_{\tilde{r}}\delta h(\tilde{r};r)=\left(\frac{2\pi}{T_{3}}\right)\frac{\delta^{(6)}(\tilde{r}-r)}{\sqrt{g(\tilde{r})}}\,.$ (83)
@@ Appendix B Late-time potential with subleading terms
$K$ $\simeq$ $-3\ln\left[\tau-\alpha\ln\tau\right.$ (86)
$+$ $\left.\frac{\xi}{3cg_{s}^{3/2}\sqrt{\tau}}\left(c-g_{s}^{2}\ln\tau\left(1-\frac{1}{3\pi\ln\tau}\right)\right)\right],$
@@
$\frac{V}{3W_{0}^{2}}\simeq\frac{\alpha}{\tau^{4}}-\frac{\xi\sqrt{g_{s}}}{4c\tau^{9/2}}\left(\ln\tau-\frac{c}{g_{s}^{2}}-b\right)+\mathcal{O}\left(\frac{1}{\tau^{5}}\right),$ (87)
with:
$b\equiv\frac{8}{3}+\frac{1}{3\pi}\simeq 2.8\,.$ (88)
@@
$\tau_{\rm min,new}=\lambda_{0}\,e^{b}\,e^{\frac{c}{g_{s}^{2}}}=e^{b}\,\tau_{\rm min}\,,$ (89)
with:
$\theta_{\rm new}\equiv\left(\frac{16c\,e^{\frac{10}{9}}}{9\xi\sqrt{g_{s}}}\right)\alpha\,e^{\frac{b}{2}}\,e^{\frac{c}{2g_{s}^{2}}}=e^{\frac{b}{2}}\,\theta\,.$ (90)
@@ C.1 Case with linear contribution
where (ignoring higher order corrections):
$\mathcal{A}$ $=$ $\frac{1}{3}+\frac{g_{s}^{2}}{3\sigma}\left(1-2\ln\sigma+3(\ln\sigma)^{2}+\frac{2\alpha}{g_{s}^{2}}\ln\sigma\right)$
$\mathcal{B}$ $=$ $-2g_{s}-\frac{2g_{s}^{3}}{\sigma}\left[1-2\ln\sigma+3(\ln\sigma)^{2}+\frac{\alpha}{g_{s}^{2}}\left(\ln\sigma+2\right)\right]$
$\mathcal{C}$ $=$ $3\left(g_{s}^{2}+\alpha\right)+\frac{3\xi}{4g_{s}^{3/2}\sqrt{\sigma}}\left[1-\frac{g_{s}^{2}}{c}\left(\ln\sigma-b\right)\right]$
$+$ $\frac{3g_{s}^{4}}{\sigma}\left[1+3(\ln\sigma)^{2}+\frac{4\alpha}{g_{s}^{2}}+\frac{1}{g_{s}^{4}}\left(2\ln\sigma\left(\alpha^{4}-g_{s}^{4}\right)+3\alpha^{2}\right)\right]$
@@
$V_{\scriptscriptstyle NNLO}$ $\simeq$ $\left(\frac{W_{X}}{W_{0}}\right)^{a}\frac{W_{0}^{2}}{\sigma^{b}}\quad\text{with}\quad a+b=5\,.$
@@ Appendix D RG-induced moduli stabilisation and inflation
In this appendix we review the RG-inspired moduli stabilisation scenario proposed in Burgess:2022nbx with the goal in mind of deepening the analysis of brane-antibrane inflation. Let us consider a string-inspired supersymmetric 4D EFT where the superpotential $W$ does not depend on the Kähler modulus $T$, i.e. $W=W_{0}$, and the Kähler potential can be expanded as:
@@
$W=W_{0}(\Phi)+XW_{X}(\Phi,\overline{\Phi})\,,$ (98)
and:
$\sigma\equiv e^{-K/3}=\tau-\hat{k}+...\,,$ (99)
@@
$\hat{k}$ $=$ $\kappa(\Phi,\overline{\Phi},\ln\tau)+(X+\overline{X})\kappa_{X}(\Phi,\overline{\Phi},\ln\tau)$ (100)
$+$ $X\overline{X}\kappa_{X\overline{X}}(\Phi,\overline{\Phi},\ln\tau)\,.$
@@
$\kappa^{\overline{X}X}\sim\mathcal{O}(1)\quad\kappa_{X\overline{T}}\sim\mathcal{O}\left(\frac{\alpha_{g,0}^{2}}{\sigma}\right)\quad\kappa_{T\overline{T}}\sim\mathcal{O}\left(\frac{k_{1}\alpha_{g,0}^{2}}{\sigma^{2}}\right),$ (102)